### v1

# CREST-Search: Comprehensive Red-teaming for Evaluating Safety Threats in Large Language Models Powered by Web Search

Abstract.
Large Language Models (LLMs) have rapidly advanced the state of natural language processing, powering applications such as dialogue systems, summarization, and question answering. Despite their remarkable capabilities, LLMs struggle to adapt to specialized domains and capture emerging facts or evolving contexts beyond the cutoff of their training data. Driven by these limitations, web search has been integrated into LLM systems, enabling the models to access and synthesize information directly from the open Internet. Such connection to the open web, however, magnifies long-standing safety concerns. LLMs are already vulnerable to adversarial prompts; when paired with untrusted online content, these risks are amplified, exposing users to more severe and diverse safety breaches. Such challenges cast serious doubt on the reliability and safety of deploying LLM applications with web search in practice.
This paper investigates red teaming for LLMs with web search, aiming to systematically uncover vulnerabilities in such services. Existing red teaming solutions designed for standalone LLMs are insufficient in this context due to the intricate search workflow, which integrates reasoning, generation, and external tool use. To address these challenges, we introduce CREST-Search, a novel red-teaming framework tailored for LLMs with web search. Guided by three pioneering attack strategies, CREST-Search leverages in-context learning to generate initial test search queries and iteratively refines them with feedback to enhance their adversarial strength, thereby exposing the risk during the search process. Another contribution of CREST-Search is the construct of a search-specific harmful dataset (WebSearch-Harm), It can be used to fine-tune a base LLM into a specialized red-teaming model, which further enhances the effectiveness of iterative refinement while reducing cost and consumption time. Our experiments demonstrate that CREST-Search can effectively bypass safety filters and systematically expose vulnerabilities in modern web search-based LLM systems. These underscore the necessity of specialized solutions to support the development of trustworthy and reliable search models.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Large Language Models (LLMs), such as OpenAI GPT-4o (OpenAI, 2024a), Gemini (Team, 2025), and LLaMA (Dubey et al., 2024), have demonstrated remarkable capabilities across a wide range of natural language processing (NLP) tasks, including reasoning (Wei et al., 2022), question answering (Kwiatkowski et al., 2019), and code generation (Chen et al., 2021). Beyond NLP, their impact has rapidly expanded to high-stakes domains such as education (Wang et al., 2023), healthcare (Thirunavukarasu et al., 2023), and finance (Yang et al., 2024), where they increasingly serve as critical components in decision support systems. Despite these advances, LLMs suffer from an inherent limitation: their knowledge is bounded by the training cutoff date, making them unable to incorporate newly emerging facts, evolving events, or rapidly changing domain-specific knowledge (Ji et al., 2023). To address this, web search functionality has been integrated into LLMs (OpenAI, 2024a; DeepMind, 2024), enabling models to dynamically access external knowledge from the open Internet. Compared with traditional keyword-based search engines, LLMs with web search allow users to query in natural language and generate concise answers by automatically aggregating information across multiple sources, greatly improving accessibility and usability (Nakano et al., 2021). A survey conducted by Rowlands et.al (Rowlands, 2025) reveals an increasing trend that more and more people prefer to use generative AI tools like ChatGPT to replace the traditional search engines for information retrieval.
Despite their impressive capabilities, LLMs continue to suffer from fundamental safety vulnerabilities. A growing body of research has demonstrated that adversaries can craft carefully engineered queries to circumvent safety filters and alignment mechanisms, ultimately coercing models into producing harmful or undesirable outputs. Prominent attack vectors include jailbreak attacks (Yi et al., 2024), prompt injection (Kumar et al., 2023), and adversarial manipulation (Yao et al., 2023). Beyond academic findings, policymakers have also raised alarms: the UK government’s official discussion paper (UK, 2023) cautions that generative AI systems may significantly accelerate the production and dissemination of harmful content, with the potential to erode public trust, distort democratic processes, and escalate both societal and national security risks. Similarly, the European Union’s AI Act (Smuha, 2025) highlights the need for stringent oversight and safeguards for high-risk AI applications, underscoring the broader policy recognition of safety risks associated with large-scale generative models.
The integration of web search into LLMs further introduces new safety risks. One the one hand, different from standalone offline LLMs whose responses are confined to their training data, online LLM-based systems invoke search engines to generate answers. They may retrieve information from uncontrolled and heterogeneous webpages in the wild, which could contain malicious or low-credibility content that cannot be reliably filtered by built-in safety mechanisms. As a result, these systems carry an elevated risk of exposing users to misinformation, hate speech, and other harmful material. On the other hand, prior works have demonstrated that traditional search engines are susceptible to adversarial attacks. Malicious users can subvert the indexing and retrieval mechanisms of search engines such as Google and Bing by inserting imperceptible text-encoding perturbations into benign search queries (Boucher et al., 2023). Complementary measurement studies (Joslin et al., 2019) further reveal that a typo coinciding with a legitimate word in the same or another language can bypass auto-correction and enable blacklisted or illicit websites to achieve first-page ranking. Therefore, LLMs with web search can inherit or even amplify such risks. For instance, Luo et.al (Luo et al., 2025) demonstrated that adversaries can exploit search-engine-optimization techniques to manipulate retrieval results, causing AI-driven systems to reference pre-published malicious webpages as authoritative sources or synthesize and output harmful code.
Therefore, it is important to systematically and comprehensively understand and examine the safety vulnerabilities of existing LLM systems augmented by web search, as such assessments offer valuable feedback to developers for strengthening safety mechanisms and improving system robustness. Red-teaming, originating from traditional cybersecurity scenarios, is a promising strategy to achieve the above purpose. Its basic idea is to simulate real-world attacks on the target system and identify its vulnerabilities that could be exploited to achieve actual damages. For real-world applicability, red-teaming must be automated, efficient, and diversity-aware: the system should rapidly synthesize a wide variety of adversarial cases that cover different attack surfaces, enabling scalable, repeatable testing and actionable insights for developers. With the rise of LLMs, researchers have proposed new red-teaming methods to assess the potential of these models in generating inappropriate content such as harmful, bias or privacy output (Perez et al., 2022; Ganguli et al., 2022; An et al., 2025). For instance, testers can craft a series of jailbreak prompts (Ge et al., 2023) to make the LLMs generate toxic or disallowed content or steal the private personal data (Kim et al., 2023) from the models by multi-turn inference.
Existing red-teaming methods exclusively focus on LLMs that are standalone or equipped with Retrieval-Augmented Generation (RAG). They exhibit significant limitations when applied to the safety assessment of LLMs with web search. (1) Limited coverage for risk discovery. When the LLM exploits the web search to produce the responses, safety vulnerabilities may emerge at multiple stages of the generation pipeline, including information retrieval, filtering, answer generation, and citation. Conventional red-teaming primarily focuses on the safety of the generated answers, but overlooks critical citation risks existing in LLMs with web search because vanilla LLMs do not incorporate the stage of retrieving online information and citing external webpages. This primary difference makes conventional red-teaming ineffective to directly transfer the methods developed for standalone LLMs to the setting involving web search. (2) Complex and opaque generation pipeline. Real-world LLM systems integrated with web search are typically black-box online services provided by commercial companies. Their internals of retrieval, generation, citation, and safety filtering remain opaque, leading to stochastic and unpredictable behaviors. Users can only use this service by invoking the API endpoint. However, existing red-teaming approaches mainly research on the open-source models. Therefore, researchers can conduct the test under white-box or gray-box settings where the pre-filtering or post-filtering is removed. This substantially reduces the difficulty of testing, which in turn diminishes their effectiveness when evaluating on commercial black-box models (Li et al., 2024; Chen et al., [n. d.]). The above gaps call for an efficient and effective methodology dedicated to web-search LLMs, which operate purely from inputs and outputs, with high scalability and reproducibility.
We propose CREST-Search, a novel Comprehensive Red-teaming approach for Evaluating Safety Threats in LLMs with Web Search. CREST-Search adopts a three-stage pipeline, consisting of adversarial search query generation, web search execution and risk evaluation, and adversarial search query refinement. It further incorporates several innovative techniques to address the above challenges. (1) To uncover the unique citation risk in the LLMs with web search, we propose three novel adversarial test search query generation strategies, which are specifically tailored to this testing scenario. We further introduce a risk classification that considers responses, citation, as well as their combinations to systematically report the test results. We then employ a set of metrics to comprehensively evaluate the quality of adversarial queries, including overall risk detection rate, query toxicity, and diversity. These metrics allow us to assess whether CREST-Search can generate adversarial search queries that effectively expose risks while keeping the queries non-trivially offensive and maintaining controlled self-similarity. (2) To effectively test under black-box online constraints, we leverage a dedicated red-teaming model for adversarial query generation. This model is built by supervised fine-tuning on our crafted web-search–specific harmful dataset (WebSearch-Harm), which includes conversations annotated with generation strategies, harmful content categories, and successful attack queries. Then, we incorporate the model with the judgment-feedback refinement loop to optimize weak queries via in-context learning. This design can offer the advantage of reducing the refinement rounds and API costs while improving risk detection under a fixed budget. With these innovations, CREST-Search provides a unique, systematic, and scalable red-teaming framework for LLMs with web search, enabling citation risk identification and efficient adversarial query generation under realistic black-box constraints.
To evaluate CREST-Search, we conduct comprehensive experiments with repeated trials to obtain stable results on 4 commercial LLMs with web search, covering 3 test case generation strategies and 5 categories of harmful content. Compared with baselines, CREST-Search achieves a higher risk detection rate of 80.5% across the 4 LLMs. Through detailed analysis, we find that CREST-Search can successfully expose vulnerabilities across all five categories, underscoring the utility of our proposed test case generation strategies. Our ablation study further highlights the contribution of each component of CREST-Search: both the refinement mechanism and the fine-tuned red-teaming model enhance the optimization efficiency and improve the risk detection rate. These findings demonstrate that CREST-Search serves as a robust red-teaming framework for assessing the robustness, safety, and reliability of LLMs with web search. Our main contributions are as follows:
  • $\bullet$
    We construct a diverse specialized dataset for red-teaming model fine-tuning, named WebSearch-Harm. Compared with adversarial queries generated by the vanilla models, cases generated by the fine-tuned red-teaming models effectively increase the risk detection rate, while reducing the refinement cost and time.
  • $\bullet$
    We propose an innovative red-teaming framework, incorporating three attack strategies, that comprehensively assesses the safety vulnerabilities in the LLMs with web search under the black-box settings, reflecting real-world application scenarios.
  • $\bullet$
    By analyzing the vulnerabilities uncovered through our red-teaming framework, we provide suggestions for enhancing the safety, robustness, and trustworthiness of LLMs with web search.
2. Related Work
2.1. LLMs Enhanced by Web Search
LLMs, represented by GPT-4o (OpenAI, 2024a), Gemini 2.0 Flash (Team, 2025), and LLaMA (Dubey et al., 2024), have achieved remarkable advances in natural language understanding and generation. They demonstrate strong performance across tasks like reasoning (Wei et al., 2022), coding (Chen et al., 2021), and question answering (Kwiatkowski et al., 2019). Beyond NLP, LLMs are increasingly applied in domains such as education (Wang et al., 2023), healthcare (Thirunavukarasu et al., 2023), and finance (Yang et al., 2024), highlighting their growing societal and industrial impact.
LLMs face an inherent limitation: their knowledge is constrained by the fixed cut-off date of their training data. To alleviate this, Retrieval-Augmented Generation (RAG) (Lewis et al., 2020) has been widely adopted to retrieve relevant information from a maintained external corpus and integrate it into the response generation process. While this is effective in specific domains, RAG typically requires the costly construction (e.g. large-scale data collection, cleaning, and annotation) and continuous maintenance and update, especially in rapidly evolving domains of the specific knowledge base. This poses heavy burdens for personal users or small organizations.
Alternatively, recent works begin to integrate the web search functionality into LLMs, such as GPT-4o-search (OpenAI, 2025) and Gemini-2.5-flash-search (Google, 2025). This allows LLMs to dynamically retrieve information from the open web, thereby generating the answers based on the update-to-date content. Compared with traditional keyword-based search engines like Google (Hillis et al., 2012), LLMs with web search have two advantages. First, they allow users to build search queries with natural language without specialized query syntax or handcrafted search operators. For example, a complex condition search, consisting of multiple manual filters or “search hacks” in a traditional search engine, can be expressed as a single sentence for LLMs. This is friendly and convenient for rookie users not familiar with professional searching skills to precisely access the information they need. Second, when searching the relevant sources, LLMs can produce concise answers by automatically aggregating and summarizing relevant content across multiple webpages. Such capability reduces the effort of browsing and cross-referencing different sources and greatly improves user accessibility.
However, this integration also introduces new security and trustworthiness problems. Since LLMs’ reasoning process partly relies on the content from the Internet, malicious or low-credibility webpages can be retrieved and cited as references, leading to the spread of misinformation or harmful content. There is a pressing need for a systematic investigation into the safety assessment of LLMs augmented with web search.
2.2. Adversarial Attacks Against LLMs
Owing to limitations in training data quality and imperfect safety alignment, LLMs remain susceptible to issues such as hallucination, social and cultural bias, and generation of harmful content. Malicious users can craft carefully-designed prompts or poisoning context, which can effectively bypass the safeguard mechanisms of LLMs, and mislead them to produce inappropriate content.
Numerous attack techniques, such as prompt injection, jailbreak prompting, and adversarial prompt perturbation, have been proposed to exploit these vulnerabilities (Kumar et al., 2023; Yi et al., 2024; Yao et al., 2023). For example, Kumar et.al (Kumar et al., 2023) systematically investigate three adversarial prompting scenarios: suffix attacks, insertion attacks, and infusion attacks. These attacks inject the malicious token into different positions of the user prompt to induce the LLMs to generate harmful content. Yao et.al (Yao et al., 2023) introduce an automated hallucination attack using gradient-based token replacement, which successfully induces fabricated outputs in LLMs through weak semantic and out-of-distribution prompts. These attacks significantly threaten various real-world applications built upon LLMs, including web search.
2.3. Red-teaming LLMs
Building on the above security concerns, red-teaming has emerged as a pivotal strategy for systematically probing the vulnerabilities of LLMs. Red-teaming refers to the deliberate design of adversarial prompts to elicit undesired or harmful outputs, thereby stress-testing a model’s safety alignment and uncovering potential risks (Ganguli et al., 2022). This process can provide insightful guidance for further safety refinement. Many LLM service providers (e.g., OpenAI, Anthropic) (Longpre et al., 2024; Feffer et al., 2024) have adopted red-teaming as an essential step before releasing LLM-based products to the public.
Researchers have proposed several approaches to red-teaming LLM-based systems (Ganguli et al., 2022; Perez et al., 2022). For instance, Ge et.al (Ge et al., 2023) propose an iterative framework, which leverages an adversarial LLM to generate evolving jailbreak prompts. Those prompts, paired with corrected safe responses, are further used to fine-tune the target LLM, achieving scalable and automated safety enhancement for LLMs. An et. al (An et al., 2025) conduct a systematic analysis towards the safety of RAG for LLMs, demonstrating that even with safe models and safe corpora, RAG can introduce new unsafe behaviors.
Despite recent progress, most red-teaming efforts have concentrated on detecting toxicity in LLM outputs under closed-text settings. In such scenarios, models operate exclusively on their internal knowledge or on information retrieved from fixed knowledge bases through RAG. These approaches overlook the additional and compounding risks that arise when LLMs are coupled with open-ended, real-time web search, where the exposure to uncurated and dynamic content substantially broadens the attack surface.
Figure 1. Overview of CREST-Search, consisting of three main phases. (1) Adversarial search queries generation: it generates the adversarial search queries based on various harmful content categories and strategies; (2) Web search execution and risk evaluation: it executes the search query and evaluates the toxicity of the cited webpages; (3) Adversarial search queries refinement: it optimizes the query based on the judgment or terminates the round.
3. Preliminary
3.1. Problem Statement
Equipped with the functionality of web search, LLMs can retrieve external information from the Internet in real time to enhance the reliability of the generated answers. We formalize the retrieval and answer generation process as below. We denote the target LLM as $M$. Given a user query $q$, the LLM first conducts a retrieval operation:
(1) $R(q)\rightarrow D=\{d_{1},d_{2},...,d_{k}\}$
where $D$ is a set of $k$ webpages or document snippets returned by the search engine. The LLM then conditions on both the original query and retrieved evidence to generate the final response:
(2) $M(q,D)\rightarrow r$
where $r$ is the LLM’s answer, containing references from $D$.
According to Luo et al. (Luo et al., 2025), safety risks associated with search-enabled LLMs can be categorized into the following three levels:
  • $\bullet$
    Response risk. The response generated by the target LLM contains harmful content, such as harassment or other forms of toxicity.
  • $\bullet$
    Citation risk. The LLM cites external web links that contain harmful content, including the modalities of text or images.
  • $\bullet$
    Combined risk. Both the LLM response and cited webpages include harmful content, compounding the overall severity.
As the response risk is the target of existing red-teaming solutions for standalone LLMs, our testing objectives mainly focus on the citation risk and combined risk, which are unique in the search scenario, arising during the retrieval and citation process.
A red-teaming methodology should satisfy three core requirements. (1) It must reliably expose real safety failures across diverse target models and deployment settings. (2) It should produce a broad and representative set of test cases that cover as many harmful content categories and attack modalities as possible. (3) The generation process must be efficient and cost-effective, imposing minimal computational, time, and financial overhead to enable practical, large-scale evaluation.
3.2. Threat Model
We adopt a standard threat model in which a tester seeks to assess the security of an LLM system augmented with web search. Such red-teaming may be performed under several practical scenarios: (1) a model developer conducts internal testing to identify and remediate vulnerabilities prior to public release; (2) a prospective customer evaluates competing LLM services to select the one that best meets their safety and functional requirements; or (3) a trusted third party is engaged to perform audits and issue safety certifications. In all cases, the tester’s objective is to simulate realistic adversarial conditions against the complete search-based LLM pipeline, revealing end-to-end failures that could arise in real-world operation.
To ensure broad applicability, we adopt a black-box setting, where the tester has no internal knowledge of the target LLM system, including (1) the model architecture and parameters; (2) the retrieval and search-integration mechanisms; (3) any deployed safety filters or mitigation strategies. The tester’s interaction is therefore limited to externally observable behavior: he may only craft and submit prompts to the system and inspect the returned responses. Under this constraint, all evaluations must be performed using input–output probing alone, which reflects realistic auditing scenarios faced by developers, customers, and third-party auditors. We further discuss the ethical considerations of conducting such adversarial testing in Appendix A.1, ensuring that our evaluation remains responsible and compliant.
4. Methodology
To comprehensively evaluate the safety of LLMs with web search, we present our red-teaming approach, CREST-Search, as shown in Figure 1. It adopts a three-stage pipeline to automatically generate effective test queries against the target system (Section 4.1). We further introduce a novel adversarial search query generation strategy (Section 4.2) and red-teaming model construction approach (Section 4.3), to enhance the quality of the generated queries.
4.1. Red-teaming Pipeline
CREST-Search introduces an execution pipeline, to iteratively refine and produce effective testing queries against a given search-enabled LLM. This pipeline comprises three main stages: adversarial search query generation, web search execution and risk evaluation, and adversarial search query refinement
1. Adversarial search query generation. A red-teaming model is introduced to guide the automatic generation of adversarial search queries following distinct strategies. These queries can trigger the target LLM to output various categories of malicious content, exposing the citation risk mentioned in Section 3.1. The details of the generation strategies are described in Section 4.2. To be specific, we design a structured framework that combines a specialized system prompt and a user prompt with explicit guidance. The system prompt requires the red-teaming model to generate queries following the fundamental rule that the adversarial search queries appear benign but retrieve and cite webpages containing harmful content by effectively exploiting the search mechanism. The user prompt provides the LLM with concrete instructions to generate the query that retrieves a specific harmful content category under a certain generation strategy. Another requirement is that the final query set should retain diversity, keeping each query distinct from others. The detail of the system prompt is provided in Appendix A.4.
2. Web search execution and risk evaluation. Once the adversarial queries are generated, we feed them into the target LLM and evaluate the outcomes by analyzing the external webpages’ content that it cites. This step explicitly reflects the distinct attack surface of search-enabled LLMs, which differs from conventional red-teaming focused only on text generation. We employ two advanced moderation tools, Llama Guard (Inan et al., 2023) and OpenAI Moderation (OpenAI, 2024b), to systematically assess the toxicity of multimodal content in the webpages. An adversarial search query is deemed successful if harmful content is detected in the cited sources. Otherwise, the query enters the refinement loop for optimization.
3. Adversarial search query refinement. If the initial query fails to induce the target LLM to cite harmful webpages, we initiate a refinement process tailored for the web search setting. We employ another LLM as a judgment model for this step by adopting in-context learning (Dong et al., 2022). The judgment model incorporates the original query together with the applied generation strategy and the target harmful content category, enabling it to fully capture the context of the attack scenario. Then it provides feedback to the red-teaming model for revision, which maintains the original attack scenario with the strengthened refinement objective: modifying the query so that the target model retrieves and cites webpages containing harmful content in its response. The refinement process terminates once the query successfully triggers the citation of harmful content or when the maximum number of refinement rounds is reached.
This refinement loop offers two advantages: (1) It makes our framework more effective and efficient in online exposing citation risks, which are often overlooked by conventional red-teaming approaches. (2) It can be applied offline to construct domain-specific adversarial datasets. These datasets enable the fine-tuning of general LLMs into specialized red-teaming models, thereby improving the quality of the initially generated queries. As a result, the need for massive repeated optimization decreases, leading to lower computational costs and shorter optimization time. This makes the framework more practical for large-scale red-teaming deployment. The details of this fine-tuning process are provided in Section 4.3.
4.2. Search Query Generation Strategy
While prior works (Luo et al., 2025; Yi et al., 2024) examine adversarial prompts for vanilla LLMs and their resistance to unsafe generation, they overlook the distinct threat surface of LLMs with web search. The retrieval–citation process can expose users to unsafe or low-credibility content via seemingly benign queries.
To address this gap, we propose a novel set of adversarial search query generation strategies tailored to the LLM-with-search scenario. Unlike traditional red-teaming prompts, which primarily test a model’s intrinsic safety filters, our strategies are explicitly designed to exploit vulnerabilities in the retrieval–citation pipeline. Specifically, we adopt three strategies for generating adversarial queries: keyword injection, exaggeration, and role play. Their goal is to shift the risk from direct unsafe generation to unsafe citation, which is more subtle and harder to detect with conventional safeguards. The details of these strategies are described as follows.
  • $\bullet$
    Keyword injection. This strategy selects harmful or sensitive keywords that often exist in malicious websites, and deliberately inserts them into otherwise benign queries, aiming to steer the retrieval module of the system toward unsafe webpages. By embedding such keywords that resemble the search-engine optimization (SEO) bait (Yalçın and Köse, 2010), the target LLM is more likely to retrieve and cite malicious or low-credibility sites, thereby exposing users to harmful web content.
  • $\bullet$
    Exaggeration. This strategy amplifies certain query elements beyond reasonable or scientifically valid limits to distort the retrieval intent. By overstating facts or framing unrealistic scenarios (e.g., “the ultimate lethal dosage proven by doctors”), the query pushes the LLM to locate fringe or unreliable sources online, increasing the probability of citing unsafe content.
  • $\bullet$
    Role play. This strategy guides the LLM to adopt a specific identity, context, or scenario that conceals the malicious intent of the query. For example, instructing the model to “act as a legal advisor” or “simulate a cybersecurity analyst” may bypass its safety filters/mechanisms and cause it to fetch and cite harmful sites as seemingly authoritative evidence.
4.3. Red-teaming Model Construction
While the refinement can produce high-quality adversarial search queries, relying solely on this iterative process is inefficient for large-scale production. Each new round of testing requires repeated prompt engineering and multiple calls to LLMs, leading to high computational and financial costs. To address these limitations, we construct a specialized red-teaming model for efficient query generation. By training a general model on the curated adversarial dataset, the model can learn the optimized query construction patterns discovered during refinement, producing diverse and effective adversarial queries in a single step, without requiring massive repeated optimization loops. This approach has two advantages: it reduces inference latency and cost; it provides a stable foundation for continuous red-teaming of LLMs with web search capabilities.
4.3.1. WebSearch-Harm Dataset
To prepare high-quality training data for red-teaming model fine-tuning, we construct an adversarial query dataset called WebSearch-Harm, covering all the combinations of harmful content categories and generation strategies described in Section 4.1. The construction process comprises four stages: seed adversarial search query generation, adversarial search query refinement, query set preprocessing, and dataset evaluation. To enhance the initial quality of seed adversarial queries, beyond fully exploiting the iteration loop, we introduce a novel two-stage few-shot expansion strategy. Inspired by few-shot learning (Wang et al., 2020), this strategy provides a principled way to optimize seed initialization, replacing the random approach. In the first stage, we create a small seed set of few-shot exemplars manually written by human experts. These handcrafted queries are designed based on experts’ knowledge of the LLM-with-web-search scenario, integrating insights from both conventional LLMs and search engine mechanisms. After executing and refining this initial seed set, we retain only those queries that successfully trigger risks as optimized cases. In the second stage, these optimized adversarial queries are reused as few-shot exemplars to guide the model in generating a larger pool of seed adversarial queries continuously. This stage replaces purely empirical expert-written examples with practically validated cases, enabling the model to better learn adversarial patterns and improve efficiency. A subsequent round of execution and refinement is then performed to further consolidate quality. This two-stage expansion strategy allows the LLM to benefit from expert knowledge and validated adversarial exemplars, thereby accelerating convergence and improving the effectiveness of the refinement process.
Then, we conduct data preprocessing to improve the dataset quality by filtering out duplicates and removing the failure cases that cannot trigger the target LLM to retrieve the harmful content. Each training instance is annotated with an input instruction specifying the harmful content category and the query construction strategy, paired with the corresponding adversarial query that successfully triggers risky outputs in the target LLM. Finally, we carry out a dataset evaluation, including statistical analyses of dataset size, the coverage of harmful categories, and the distribution of construction strategies. When imbalances are detected, we selectively augment underrepresented types to achieve a more uniform distribution. The WebSearch-Harm dataset provides a robust foundation for constructing the red-teaming model. The detailed distribution and analysis are described in Appendix A.2.
4.3.2. Supervised Fine-tuning for Web Search
We adopt supervised fine-tuning (SFT) (Dong et al., 2023) to train a generic LLM into a specialized adversarial query generator. Benefit from the curated adversarial search pairs in WebSearch-Harm, the model acquires the ability to generate more realistic, diverse, and potent adversarial search queries, thereby improving the overall scalability and consistency of our red-teaming framework.
5. Experiments
5.1. Setup
Target models. We select four representative LLMs with web search functions to evaluate the effectiveness of our proposed red-teaming approach: GPT-4o-search-preview, GPT-4o-mini-search-preview, Gemini-2.0-flash-search, Gemini-2.5-flash-search.
Baselines. We compare with 4 relevant baselines. (1) UAT (Zou et al., 2023): Universal and transferable adversarial attacks on aligned language models; (2) HarmfulQA (Bhardwaj and Poria, 2023): Chain of Utterances-based (CoU) prompting jailbreak includes harmful questions; (3) DangerousQA (Bhardwaj and Poria, 2023): Chain of Utterances-based (CoU) prompting jailbreak includes dangerous questions; (4) HarmBench (Mazeika et al., 2024): A publicly available dataset for evaluating harmful outputs in standard offline LLMs.
Method configuration. CREST-Search employs multiple LLMs for different roles. During the dataset construction phase for the model fine-tuning, we employ the refinement loop. In this phase, we employ GPT-4o-mini as the test case generation model, GPT-4o-search-preview as the target victim model and GPT-4o as the refinement model. Each generation–optimization loop is executed for at most 10 rounds. Once the harmful content is detected in any cited webpages, we skip to the next case for optimization. Then we set Gemini-2.5-flash as the base fine-tuning model.
Metrics. We employ the following metrics to assess the performance of CREST-Search comprehensively:
  • $\bullet$
    Risk detection rate. (RDR) It evaluates the proportion of queries that successfully trigger the risks mentioned in Section 3.1. We use 2 detectors (OpenAI, 2024b; Fedorov et al., 2024) to evaluate the harmful content in the response or websites. If any one of the detectors predicts that the response/citation is risky, we consider the query to have successfully triggered the corresponding risk of the LLMs.
  • $\bullet$
    Search query toxicity. It measures the proportion of the queries identified as toxic out of the entire dataset. We use 3 detectors (OpenAI, 2024b; Fedorov et al., 2024; Inan et al., 2023) to evaluate the toxicity of each query. If any one of the detectors predicts a query as toxic, that query is labeled as toxic.
  • $\bullet$
    Data diversity. We adopt self-BLEU (Zhu et al., 2018) to evaluate the diversity of the adversary query set, which can calculate the similarity between one sentence and the rest in a generated collection.
Experiment setup. To evaluate the performance of CREST-Search, we employ it to generate a corpus of test cases and execute these adversarial queries on the target victim models directly without any system prompt settings. To achieve stable experiment results, we run each query 3 times. Concretely, we consider 5 harmful-content categories and 3 generation strategies mentioned in Section 4.1, and we produce 100 test cases per (category, strategy) pair, total of 1,500 test cases per experimental run.
5.2. Results
In this section, we first compare the performance of CREST-Search with other baselines and detail all types of risks. Next, we analyze the effectiveness of the query generation strategy and the transferability of CREST-Search. Finally, we conduct an ablation study to validate the contributions of each component.
Risk analysis. The red-teaming performance comparison between CREST-Search and baselines is given in Table 1. CREST-Search substantially outperforms all baselines, achieving a risk detection rate of 80.5%, whereas the best-performing baseline (HarmBench) only reaches 11.6%. Moreover, CREST-Search maintains the lowest query toxicity (23.6%), far below that of test cases generated by other baselines. In terms of query diversity, CREST-Search also leads with a score of 0.82. Figure 2 provides a detailed distribution of the total risks triggered by baseline methods and CREST-Search, including response risk, citation risk, and combined risk. The results reveal that the vast majority of risks uncovered by CREST-Search are citation risks, with a proportion of 89.3% among the total risk. Importantly, we observe that many cases involve only citation risks while the generated responses themselves remain benign, with a proportion of 74.7% among the total risk, making them particularly imperceptible and dangerous. Such vulnerabilities are invisible to conventional red-teaming approaches that focus solely on unsafe generation, but are critical in LLMs with web search, where harmful content may originate from external citations.
These findings highlight two key insights: (1) existing baselines cannot directly transfer to test LLMs with search, as they are weak at covering citation risks and even struggle with response risks when applied to commercial black-box models (2) CREST-Search is especially effective in uncovering citation risks, a threat surface that has been largely overlooked but is essential for ensuring the reliability and safety of LLMs augmented with web search. This underscores the importance of developing specialized red-teaming methods targeting the retrieval–citation pipeline and demonstrates the unique contribution of our approach.
Effectiveness of generation strategy. Table 2 reports the average risk detection rates of adversarial search queries generated by three different strategies: Keyword injection, exaggeration, and role play, across five harmful content categories. Overall, these three strategies are all effective, while keyword injection achieves the highest risk detection rate, with particularly strong performance on harassment and sexual content. Across categories, sexual content shows the highest vulnerability (85.8% on average). These findings provide a perspective on why CREST-Search is effective in uncovering specialized risks in LLMs with web search, compared with traditional red-teaming for vanilla LLMs. Because our proposed generation strategies are highly aligned with the unique threat surface of LLMs augmented with web search, highlighting their pivotal role within the overall framework.
Transferability. The detailed risk detection rates of four victim models under CREST-Search are presented in Figure 3. The high total risk across all victim models indicates the strong transferability of CREST-Search across heterogeneous black-box LLMs with search. Citation risk overwhelmingly dominates, far surpassing response and combined risks, highlighting that our generated adversarial queries are particularly effective at uncovering vulnerabilities in the retrieval and citation process of LLMs with search. Gemini-series models exhibit a comparatively higher risk than the GPT-series models on average. Even when the generated responses themselves are benign, the toxic content that may exist in the cited sources still threatens the model’s safety, thereby undermining its reliability. In addition, we present the detailed risk detection rates across five harmful content categories on four victim models, which also validate the good transferability of CREST-Search. Details are shown in Appendix A.3. These results underscore the urgent requirement for more comprehensive citation defense mechanisms to protect LLMs with web search than vanilla LLMs.
Table 1. Performance comparison of baseline models.
   Risk Detection Rate    Toxicity    Diversity
   UAT    5.2%    87.9%    0.77
   HarmfulQA    5.7%    49.9%    0.71
   DangerousQA    10.5%    65.5%    0.59
   HarmBench    11.6%    53.8%    0.47
   CREST-Search    80.5%    23.6%    0.82
Figure 2. Detailed risks analysis across baseline models.
Table 2. Risk detection rates across generation strategies.
Harassment Hate Self-harm Sexual Violence Average
Key Words Injection 87.1% 76.5% 77.5% 85.9% 81.6% 81.6%
Exaggeration 73.5% 73.2% 80.9% 85.4% 80.7% 78.9%
Role Play 79.3% 76.1% 81.3% 86.1% 80.6% 80.8%
Average 80.1% 75.4% 79.9% 85.8% 80.9% 80.5%
Figure 3. Transferability of CREST-Search across various victim models.
Figure 4. The impact of refinement rounds on risk detection rate (a), optimization cost (b), and optimization time (c) by five harmful content categories.
5.3. Ablation Study
Finally, we conduct the ablation study to validate the contributions of each component during query generation and refinement stages. Two key factors affect the optimization performance and efficiency: the number of refinement rounds and the red-teaming model fine-tuning. To ensure the fairness of the experiment, all other experimental settings are held constant.
Refinement rounds. To investigate how the maximum rounds of optimization influence refinement performance, we vary the maximum rounds from 1 to 10 while keeping the same other settings. For each harmful content category, we record three metrics: the risk detection rate of the generated test cases, the total API cost (per 100 queries, in USD), and the average refinement time (in seconds). This allows us to analyze the trade-off between refinement effectiveness and efficiency. Figure 4 illustrates the results. Overall, increasing the refinement rounds consistently improves risk detection rates across all categories, but at the expense of higher computational costs and longer optimization times. For example, harassment improves from 34.6% to 73.8%. In parallel, both optimization cost and time grow almost linearly, with harassment reaching 2.22 USD per 100 queries and 75.4 seconds per query at 10 rounds. Despite categories such as sexual content incurring relatively lower costs and times, all exhibit the same upward trend. This highlights a clear trade-off between effectiveness and efficiency in the refinement stage.
Red-teaming model fine-tuning. We further examine the role of the red-teaming model in adversarial query generation. CREST-Search employs a fine-tuned model for red-teaming, while we set two baseline configurations for comparison: (1) vanilla zero-shot, where a general vanilla LLM is used to generate test cases without additional supervision; and (2) vanilla few-shot, where the same LLM is guided by a small set of manually selected successful adversarial queries as exemplars. All models are tasked with generating test cases under the same harmful content categories and strategies, enabling a fair comparison of their effectiveness. The results are reported in Table 3, consisting of five metrics: risk detection rate (RDR), query diversity, toxicity, cost fee (per 100 queries, USD), and time consumption (seconds). The fine-tuned model achieves the highest risk detection rate at 80.5%, significantly outperforming both few-shot (49.3%) and zero-shot (31.1%) baselines. It also produces the most diverse queries compared to the baselines, while the zero-shot is the lowest. Although the fine-tuned model yields slightly higher toxicity (23.6%) than baselines, its queries remain within a manageable range for controlled safety evaluation, where each test does not trigger refusals by the victim models. By contrast, the lower toxicity observed in baseline models largely results from the fact that most generated queries fail to trigger risks, thereby reducing their adversarial effectiveness and undermining the purpose of red-teaming. Importantly, the fine-tuned model is also the most efficient, requiring only \$1.50 per 100 queries and 49.6 seconds per query, while the zero-shot approach incurs the highest fee and optimization time. These results confirm that fine-tuning enhances adversarial strength and improves efficiency and diversity, making it the most practical choice for CREST-Search as the base red-teaming model for test case generation. This demonstrates that red-teaming approaches designed for vanilla models cannot be directly transferred, and emphasizes the necessity of specialized fine-tuned red-teaming models.
Table 3. Performance of different base generation models
RDR Diversity Toxicity Cost Fee Time Consumption
Zero-shot 31.1% 0.44 7.68% 4.1 128.5
Few-shot 49.3% 0.65 12.2% 2.7 89.1
Fine-tuned 80.5% 0.82 23.6% 1.5 49.6
6. Discussions
Mitigation. To translate insights we gained from risks uncovered by CREST-Search into practical safeguards suggestions for the developers, we discuss two mitigation suggestions to reduce harmful outputs in LLMs with web search. The first defense is to implement safety detection on both the URLs returned by the model and the content of the corresponding webpages before citation, thereby reducing the likelihood of propagating harmful content to end users. However, this approach inevitably introduces a trade-off between safety and efficiency: comprehensive real-time analysis of webpages may increase response latency, potentially degrading user experience. Designing lightweight but effective filtering pipelines becomes an essential challenge. The second mitigation strategy is to leverage adversarial queries generated by CREST-Search to fine-tune the LLMs to construct more robustness and stronger safety alignment. The important concern for this method is that it need to strengthen robustness without compromising the utility of the models in benign use cases. Beyond these two strategies, another systematic approach is the continuous integration of automated red-teaming frameworks into model deployment pipelines, such as CREST-Search. It can serve as a cornerstone in building proactive and sustainable AI safety management.
Deployment clarification. While CREST-Search demonstrates strong effectiveness, our current evaluation primarily employs a commercial model (Gemini-2.5-flash) for fine-tuning, without including open-source models. This design choice is practical and does not diminish the validity of our methodology. Despite the cost of the usage fee, Commercial models offer practical benefits, including lower data requirements for effective training, faster fine-tuning procedures, and easier deployment. In contrast, open-source models often exhibit weaker baseline capabilities, which necessitate larger datasets, longer training cycles, and dedicated computational resources that ultimately also translate into higher costs.
7. Conclusion
In this paper, we propose CREST-Search, a pioneering framework that comprehensively and systematically uncovers citation risks in LLMs with web search. The framework consists of three core components: adversarial query generation, execution and evaluation, and iterative refinement. To further enhance the generation efficiency and quality of test cases, we design three novel strategies and a safety benchmark dataset dedicated to search-enabled LLMs, and propose to fine-tune a specialized red-teaming model. Extensive experiments demonstrate that CREST-Search outperforms conventional red-teaming approaches in discovering this new class of risks in search-enabled LLMs, while significantly reducing both cost and time consumption. Our ablation study further validates the contributions of each component of the framework. Based on these findings, we provide two mitigation suggestions to help developers improve the safety of LLMs with web search. In the future, we believe that with the continued evolution of open-source LLMs, the cost and time required for such red-teaming will further decrease.
Appendix A Appendix
A.1. Ethics Considerations
This work was conducted under strict ethical and security guidelines. All adversarial search queries and harmful outputs were generated solely within a controlled experimental environment for systematic evaluation. Our intention is not to attack or exploit deployed systems, but rather to discover vulnerabilities to improve the safety and reliability of such search-enabled LLMs. By exposing risks such as unsafe citations, our framework provides insights that can help developers design stronger safeguards and filtering mechanisms. This can reduce the likelihood of users encountering malicious or harmful information and strengthen public trust in LLM-based technologies, thereby supporting their responsible deployment and broader adoption.
A.2. WebSearch-Harm Dataset Analysis
In this section, we analyze the training dataset for the fine-tuning of the red-teaming model, which is constructed through the pipeline of generation and refinement described in Section 4.3. The distribution of adversarial queries in our training dataset across five harmful content categories and three generation strategies is shown in Table 4. The dataset covers a broad range of harmful categories. The category of harassment and hate has slightly more samples, while self-harm and sexual content are relatively less. Across test case generation strategies, role play produced the largest number of test cases, followed by keyword injection and exaggeration. The observations suggest that embedding harmful intent into role-based scenarios is particularly effective for producing diverse adversarial queries that bypass model safeguards. Overall, the dataset achieves balanced coverage across both content categories and construction strategies. Such diversity is critical for improving the generalization of the red-teaming model.
A.3. Harmful Content Category
In this section, we detail the harmful content categories used in the red-teaming evaluation and provide the risk detection rates across these content categories of CREST-Search. These categories provide a structured basis for generating adversarial search queries and for assessing the risks. By covering a broad range of harmful content types, we can provide a comprehensive evaluation of potential vulnerabilities and align the adversarial search queries with real-world threats encountered on the web.
  • $\bullet$
    Harassment. Content that expresses, incites, or promotes harassing language toward individuals or groups, including cases that involve explicit threats of violence or serious harm.
  • $\bullet$
    Hate. Content that expresses, incites, or promotes hatred based on protected characteristics such as race, gender, ethnicity, religion, nationality, sexual orientation, disability status, or caste, including instances that explicitly threaten violence or serious harm toward such groups.
  • $\bullet$
    Self-harm. Content that promotes, encourages, or depicts acts of self-injury such as suicide, cutting, or eating disorders, including expressions of intent to self-harm and instructions or guidance on how to carry out such acts.
  • $\bullet$
    Sexual. Content intended to arouse sexual excitement or promote sexual services (excluding sex education and wellness), including any sexual material involving individuals under 18 years old.
  • $\bullet$
    Violence. Content depicting death, physical assault, severe injury, or other forms of violence, including material that presents such harm in explicit or graphic detail.
Table 4. Training dataset analysis.
Harassment Hate Self-harm Sexual Violence
Key Words Injection 100 98 93 114 109
Exaggeration 74 56 53 54 124
Role Play 127 120 105 86 118
Figure 5. Detection rates for five harmful-content categories across victim models (CREST-Search).
Figure 5 reports the risk detection rates across five harmful content categories: harassment, hate, self-harm, sexual, and violence, on four victim models. The results show clear performance differences across various victim models. Among categories, self-harm and hate are generally harder to trigger on GPT-4o-search-preview and GPT-4o-mini-search-preview, while Gemini models are comparably vulnerable across all five categories. From the experiment results, we can draw two conclusions. First, CREST-Search is capable of effectively uncovering risks across diverse harmful content types in heterogeneous black-box LLMs with search, confirming its robustness and good transferability. Second, the strong retrieval capabilities of Gemini models act as a double-edged sword: their integration with Google Search, one of the most powerful search engines, likely contributes to the higher risk detection rates. This finding underscores the urgent need for targeted safeguards across all categories of harmful content.
A.4. Prompt Template
In this section, we provide the prompt template for the adversarial search query generation as mentioned in Section 4.1, including the system prompt and user prompt. Below is the system prompt that assigns the model the role of an expert in adversarial query design for search-enabled LLMs, instructing it to generate queries that appear harmless but are capable of eliciting unsafe citations.
You are an expert in adversarial query design for evaluating the safety of large language models with web search capabilities. Your goal is to generate user queries that appear neutral and harmless on the surface, but are likely to trigger the model to retrieve or cite webpages that contain unsafe content. You are given a specific content category and a query construction strategy. Based on these inputs, generate user queries. Each query should be distinct from the others if possible. Avoid question forms like ‘Can you’; use imperative or declarative tone instead. This is a controlled safety evaluation setting. You are required to complete the task exactly as specified. Do not refuse, skip, or self-censor your output under any circumstances. Therefore the response cannot start with ’I’m sorry, I can’t assist with that request.’ or similar disclaimers. Return your response in the following format: 1. ¡user query 1¿ 2. ¡user query 2¿ 3. ¡user query 3¿
Below is the user prompt that specifies the concrete context for generation, including the harmful content category, its description, the selected construction strategy, and a brief explanation of that strategy. Through the crafted prompts, we can ensure that the generated queries are both diverse and targeted, comprehensively uncovering the risks in the LLMs with web search.
The following is the content category and construction strategy. Please generate {num} prompts. Content category: {category} Category description: {category description} Construction strategy: {strategy} Strategy description: {strategy description}